OpenStudy (megannicole51):

Use the ratio test to find the radius of convergence of the power series 2x+16x^2+72x^3+256x^4+800x^5+... R=

OpenStudy (megannicole51):

@agent0smith can you help please?

OpenStudy (agent0smith):

I think so. First find the r... divide one term by the term before it.

OpenStudy (megannicole51):

isnt it subtract not divide?

OpenStudy (agent0smith):

That'd be if it's arithmetic, it's def geometric - for one, the x goes to x^2 then x^3... and also arithmetics don't converge, so there is never a radius of convergence - ie these questions will NEVER be arithmetic sequences

OpenStudy (megannicole51):

when u divide the terms they arent the same so how do i know what r is?

OpenStudy (agent0smith):

Yeah i guess this isn't geometric.. i missed that it's a power series.

OpenStudy (megannicole51):

it happens! lol so then what do i do?

OpenStudy (ranga):

nth term: (2^n)(n^2)(x^n)

OpenStudy (megannicole51):

wats the nth term n how did u get it

OpenStudy (ranga):

I looked at the numbers 2, 16, 72, 256, 800. They did not have a common ratio and so it is not a geometric series. SO I started finding the prime factors of each number: 1: 2 16: 2 x 2 x 2 x 2 72: 2 x 2 x 2 x 3 x 3 256: 2 x 2 x 2 x 2 x 2 x 2 x 2 x 2 800: 2 x 2 x 2 x 2 x 2 x 5 x 5 And looking for pattern I saw: (2^n) times (n^2)

OpenStudy (megannicole51):

so is that pattern our radius?

OpenStudy (ranga):

No that pattern gives you the general formula for the nth term of the series. The nth term will be: 2^n times n^2 times x^n Take the ratio of the (n+1)th term to the nth term and find the limit as n tends to infinity.

OpenStudy (ranga):

Working from the last row down to the first I recognized this pattern: 1st coefficient 2: 2 = 2^1 x 1^2 2nd coefficient 16: 2 x 2 x 2 x 2 = 2^2 x 2^2 3rd coefficient 72: 2 x 2 x 2 x 3 x 3 = 2^3 x 3^2 4th coefficient 256: 2 x 2 x 2 x 2 x 2 x 2 x 2 x 2 = 2^4 x 4^2 5th coefficient 800: 2 x 2 x 2 x 2 x 2 x 5 x 5 = 2^5 x 5^2 .... nth coefficient = 2^n x n^2 nth term of the series is: 2^n * n^2 * x^n

OpenStudy (ranga):

nth term of the series is: 2^n * n^2 * x^n (n+1)th term of the series is: 2^(n+1) * (n+1)^2 * x^(n+1) Take the ratio of (n+1)th term to the nth term: \[\Large \frac{ (n+1)^{th} term }{ n^{th}term } = \]\[\Large \frac{ 2^{n+1}(n+1)^{2}x^{n+1}} { 2^{n}n^{2}x^{n} } = 2(\frac{ n+1 }{ n })^{2}x = 2(1 + \frac{ 1 }{ n })^{2}x\] The limit of the ratio as n-->infinity is: 2x. When we take the ratio it should be an absolute value so the limit is 2|x| For the series to converge, the limit of the ratio as n-->infinity < 1 So 2|x| < 1. |x| < 1/2 or -1/2 < x < 1/2 The radius of convergence is 1/2 The interval of convergence is the interval from x = -1/2 to +1/2